OpenStudy (anonymous):

A drama club is planning a bus trip to New York City to see a Broadway play. The table represents the cost per person for the bus rental compared to the number of people going on the trip. What function models the data, and how much per person will it cost if 12 students go on the trip? Number of Students(n) - Cost per Student(c) 3 - 24$ 6 - 12$ 9 - 8$ 16 - $4.5 A. n/c = 72, $12 B. nc = 9, $10 C. nc = 72, $6 D. n/c = 9, $12

hartnn (hartnn):

As u can see the product of number of students(n) and cost per student is(c) always constant = 72 3*24=72 12*6=72 and so on. so nc = 72 now if n=12, c= 72/12 = ??

hartnn (hartnn):

@alyssababy7 did u get what i have done ?

OpenStudy (anonymous):

yes thank you!

hartnn (hartnn):

welcome :)